OpenStudy (sachin00007):

A gun of 5 kg shot out a bullet of 5g at 100 m/s what is the recoil velocity of the gun?

OpenStudy (sachin00007):

yes

OpenStudy (sachin00007):

linear momentum =0.5

OpenStudy (anonymous):

the recoil velocity of the gun is 0.04 m/s.

OpenStudy (anonymous):

explanation m1=5kg=5000g u1=0 v1=x m/s m2=5g u2=100 m/s v2= 0 m1u1+m2u2=m1v1+m2v2 5000*0+5*100=5000*x+5*0 500=5000x therefore, x=500/5000=0.04 m/s.